A square tile has a side of 3 cm. If a patio of 900 cm2 has to be tiled how many of these tiles will be needed?
Answer by amfagge92(93)

Area of a tile =3*3=9cm2
area of the patio=900cm2
number of tiles=900/9=100tiles.